# DualMap: Online Open-Vocabulary Semantic Mapping for Natural Language Navigation in Dynamic Changing Scenes

### v2

@@ I Introduction
Imagine a robot navigating a bustling office when a colleague suddenly requests, “Bring me the MacBook I was using earlier.” This seemingly simple task requires a mapping system that bridges human language and dynamic environments. To succeed, three critical capabilities emerge: 1) Open-vocabulary: The system must detect objects in a class-agnostic manner during mapping and interpret various natural language queries (e.g. ‘the MacBook’) without relying on a set of fixed categories; 2) Efficient online mapping: The system must incrementally construct and maintain semantic maps in real-time, allowing it to adapt to ongoing changes in the environment; 3) Navigation with dynamic changes: In human-centered environments, objects frequently change their location or state—for instance, a laptop might be moved or opened. The system should implement a navigation strategy that guides the agent to the relocated object and updates the map for lifelong navigation.
@@ II-A Open-Vocabulary Online Semantic Mapping
Fast closed-set detectors like YOLO [14] have been widely used in semantic robotic mapping [15, 16, 17, 1], but their fixed category sets limit generalization to open-world settings. To enable open-vocabulary perception, two main approaches have emerged. One uses image tagging models [18, 19] to generate class labels from incoming frames, which then guide closed-set detectors [20] for object detection. ConceptGraphs [4] follows this pipeline but remains offline due to the high cost of running GroundedSAM [21] per frame. A more common alternative applies class-agnostic segmentation [7, 8] directly, followed by feature embedding using vision foundation models (VFMs) [22, 23, 24, 25]. HOV-SG [5] uses this method and merges over-segmented objects at the 3D level, improving semantic accuracy but remaining expensive in computation. CLIO [26] addresses efficiency by merging only task-relevant objects. However, since the map is task-specific, querying non-task objects requires extra map reconstruction, limiting map reusability. Different from prior approaches, DualMap adopts a hybrid segmentation strategy for efficient, holistic open-vocabulary mapping and replaces costly 3D inter-object merging with lightweight intra-object checks, enabling online operation.
@@ III-A4 Scene Pointcloud Generation
To support layout estimation for downstream navigation, we maintain a scene point cloud $\mathbf{P}_{\text{s}}$ that captures the overall geometry information of the environment during operation (Fig. 3-c). It is built by selecting RGBD frames with significant pose changes (translation $>$ 1.0m or rotation $>$ 20°), projecting their point clouds, and applying downsampling. This process (blue branch in Fig. 3) runs in a parallel low-frequency thread, providing a compact structural representation of the scene.
Figure 3: Observation generation pipeline. From each posed RGBD frame, the system runs three parallel threads: a) Closed-set detection identifies in-list objects in object-level but misses unseen ones (e.g., stools); b) Open-set segmentation captures a broader object range but may over-segment (e.g., window). After refinement and merging, the results are embedded into the final observation set. c) Scene cloud generation runs at low frequency and generates 3D point clouds for layout calculation.
@@ IV-A Map Abstraction
The abstraction of the concrete map involves two steps: first, identifying which objects are selected as anchor objects; and second, associating the remaining volatile objects with anchor objects based on spatial relations. Anchor objects refer to static items that rarely move, such as beds or desks, while volatile objects (e.g., books or cups) frequently change positions. Only anchor objects are retained in the abstract map. To identify them, we construct a list of static categories and descriptive phrases (e.g., “furniture that is not often moved”), encode them with CLIP, and average the embeddings to obtain a template feature $\mathbf{f}_{\text{t}}$. For each object $o\in\mathcal{M}_{c}$, we compute the cosine similarity between an object’s feature $\mathbf{f}_{o}$ and the anchor template feature $\mathbf{f}_{\text{t}}$. If the similarity exceeds a threshold $\tau_{a}$, the object is classified as an anchor and denoted as $a$; otherwise, it is considered volatile and denoted as $v$.
@@
Scene layout, important for navigation, is also extracted in map abstraction. We project the scene point cloud $\mathbf{P}_{\text{s}}$ onto a bird’s-eye-view plane and discretize it into spatial bins. A binary occupancy map is then generated by computing a histogram of point counts across bins and using the 90th percentile as a threshold, highlighting scene structures like walls. The layout is visualized as gray point clouds in Fig. 1.
@@ IV-C Navigation Strategy
A global path toward ${a^{*}}$ is planned using a Voronoi-based planner over the abstract map. As the agent moves along this path, it incrementally builds a local concrete map $\mathcal{M}_{c}^{\text{local}}$ and evaluates objects in $\mathcal{M}_{c}^{\text{local}}$ for potential matches. For each object $o\in\mathcal{M}_{c}^{\text{local}}$, we compute the cosine similarity $s(o)$ between its feature $\mathbf{f}_{o}$ and the query feature $\mathbf{f}_{q}$. If $s(o)$ is within a margin $\epsilon$ of $s(a^{*})$ and $o$ lies within the projected footprint of $a^{*}$, it is considered a confident match. A local path is then planned via RRT* to reach the target.
@@ V Experiments
This section first describe the experimental setup and further evaluates DualMap across four core aspects: 1) the effectiveness and efficiency of online concrete map construction; 2) the navigation performance in both static and dynamic scenes; 3) the contribution of key system components (via ablations); and 4) real-world applicability.
@@ V-A1 Datasets
We evaluate DualMap on semantic segmentation, efficiency, and object navigation. For semantic segmentation and efficiency, we use the ScanNet [12] and Replica [11] datasets, following scene selections from prior work [4, 5]. For static object navigation, we evaluate on the HM3D dataset [13], selecting three representative scenes (00829, 00848, 00880) and augmenting object diversity by integrating YCB objects [32]. For dynamic navigation, as current benchmarks lack object-level dynamics, we construct new environments for dynamic object navigation using a custom tool built on Habitat Simulator [33], which enables dynamic placement of YCB objects. We define two types of object changes: In-anchor relocation, where an object moves within the same anchor (e.g., moving a cup on a table); Cross-anchor relocation, where an object is moved to a different anchor (e.g., from a table to a shelf). For each static HM3D scene, we create dynamic variants by manually repositioning added objects to reflect both types of changes. Agents are deployed in the Habitat Simulator for both static and dynamic navigation experiments.
V-A2 Evaluation Metrics
For semantic segmentation, we adopt standard metrics from [5], including mean IoU (mIoU), frequency-weighted IoU (F-mIoU), and mean accuracy (mAcc). We also report the Object Density Ratio (ODR), defined as the ratio of predicted object counts to ground-truth object counts. This metric captures how well the system reproduces realistic object quantities; values closer to 1 indicate better alignment with real-world object density. For efficiency, we measure average memory usage, peak memory consumption, and time per frame (TPF) during operation. For navigation, we report Success Rate (SR), defined as the percentage of queries in which the agent stops within 1 meter of the queried object. In dynamic scenes, success further requires finding the target within three attempts.
@@ V-A2 Evaluation Metrics
TABLE I: Open-vocabulary 3D Semantic Segmentation and Efficiency
Replica
Method mIoU $\uparrow$ FmIoU $\uparrow$ mAcc $\uparrow$ ODR $\approx 1$ Avg. Mem (MB) $\downarrow$ Peak Mem (MB) $\downarrow$ TPF (s) $\downarrow$
ConceptGraphs 0.1501 0.3858 0.3559 2.02 7148.9 23551.9 4.188
@@
ScanNet
Method mIoU $\uparrow$ FmIoU $\uparrow$ mAcc $\uparrow$ ODR $\approx 1$ Avg. Mem (MB) $\downarrow$ Peak Mem (MB) $\downarrow$ TPF (s) $\downarrow$
ConceptGraphs 0.0882 0.3077 0.3538 6.97 9780.3 26155.2 6.301
@@ V-A4 Implementation Details
All experiments are conducted on a machine with an NVIDIA RTX 4090 GPU and an Intel i7-12700KF CPU. We adopt the YOLOv8l-world model [20] for closed-set detection, using a class list generated by GPT-4 [31], and supplement it with FastSAM-s [8] to enable open-vocabulary segmentation. For feature embedding, we use MobileCLIP-S2 [25], shared across all methods for fair comparison. For real-world tests, we deploy a wheeled robot platform equipped with an RGBD camera and a LiDAR for pose estimation, with all modules communicating with DualMap via ROS2.
@@ V-B Concrete Mapping Evaluation
We present quantitative and qualitative results in Table I and Fig. 5. DualMap achieves the best semantic segmentation results across datasets, improving mIoU on Replica by 2.8% over HOV-SG and 10.3% over ConceptGraphs. This gain stems from object-level splitting that preserves objects’ diversity and text-enhanced feature embedding. Fig. 5 further illustrates these improvements, with red arrows showing misclassified objects that are better handled by DualMap. For ODR, DualMap achieves the closest alignment with ground truth (0.97 on Replica), thanks to holistic segmentation that reduces over-segmentation and our stability check that removes noisy objects. While HOV-SG and ConceptGraphs apply inter-object merging to control segment counts, they still suffer from redundant fragments. This difference is clearly illustrated in Fig. 5, where noisy RGB-D inputs from real sensors result in meaningless predictions (blue arrows) in the baselines. DualMap filters out such errors via stability check, demonstrating robustness in real-world scenarios.
@@ V-C Navigation in Simulated Environments
TABLE II: Object Navigation Success Rate (SR) on HM3D.
@@
For static navigation, we evaluate all methods across three HM3D scenes. As shown in Table II-Static, DualMap achieves the highest success rate. ConceptGraphs performs competitively due to its object-centric mapping. In contrast, HOV-SG’s segment-merging strategy still results in fragmented maps, leading to more query and navigation failures.
For dynamic scenes (Table II-Dynamic), DualMap naturally handles in-anchor changes, as the queried object remains within its original anchor and can be accurately localized through the local concrete map. For more challenging cross-anchor cases, it updates the abstract map online and reselects reasonable candidates accordingly. This strategy maintains high success rates even under large object movements. Among failure cases in cross-anchor setting, 28.3% are due to false matches, 9.4% to exceeding navigation attempt limits, and 1.9% to planning errors.
@@ V-D1 Mapping Performance
TABLE III: Semantic Mapping Ablation on Replica
@@
Full System 0.5207 0.4024 0.2537
In Table III and Fig. 5, we ablate key components of the concrete mapping pipeline on Replica and ScanNet. We first remove FastSAM and rely only on YOLO for detection. Since YOLO detects only predefined categories, many unseen objects are missed, leading to a drop in performance. Next, we disable the refinement step that merges overlapping YOLO detections based on color histogram similarity. Without merging overlaps, objects are over-segmented. The performance drops slightly, as YOLO produces relatively holistic detections. Then, we remove the weighted feature merging and use only image features. This causes a notable drop, indicating the importance of incorporating global semantic cues from text. Finally, we disable the object-level splitting mechanism. This causes under-segmentation, where nearby objects are grouped as one, leading to inaccurate matching and reduced mapping precision. Qualitative results in Fig. 5 show that without object stability check, sensor noise introduces noisy segments, degrading overall semantic quality.
@@ V-D2 Efficiency
TABLE IV: Mapping Efficiency Ablation on ScanNet. Gray row indicates configuration used in our purposed system.
@@
In Table IV, we analyze three key modules that are related to our mapping pipeline to evaluate their impact on efficiency: Inter-object merging in 3D, intra-object level stability check, and closed-set detection assistance in open-vocabulary segmentation. HOV-SG follows the configuration in Row 1, resulting in the highest time cost and ODR. This is because SAM-style segmentation produces a large number of segments, making 3D-level merging computationally expensive. Configurations in Row 2 improve speed by using YOLO to assist segmentation, but still suffer from noisy objects caused by sensor noise. In Row 3, enabling our proposed stability check significantly improves ODR by removing noisy objects, which also reduces the object merging time cost. Our final configuration (Row 4) disables inter-object merging, slightly sacrificing map accuracy but achieving notable efficiency gains—an important trade-off for real-world robotic applications.
@@ V-D3 Dynamic Scene Navigation
TABLE V: Success Rates under Different Candidate Selection Strategies for Relocated Objects on HM3D
@@
To evaluate navigation under dynamic changes, we focus on the more challenging cross-anchor setting and test different candidate selection strategies when the queried object has changed position. As shown in Table V, random selection often fails due to excessive retries. Using the original abstract map $\mathcal{M}_{a}$ provides anchor-based guidance and improves success. Our final strategy updates $\mathcal{M}_{a}$ during navigation to form $\mathcal{M}_{a}^{\prime}$, which makes full use of past navigation information. This helps the agent find the relocated object more effectively.

### v3

@@ I Introduction
Imagine a robot navigating a bustling office when a colleague suddenly requests, “Bring me the MacBook I was using earlier.” This seemingly simple task requires a mapping system that bridges human language and dynamic environments. To succeed, three critical capabilities emerge: 1) Open-vocabulary: The system must detect objects in a class-agnostic manner during mapping and interpret various natural language queries (e.g., ‘the MacBook’) without relying on a set of fixed categories; 2) Efficient online mapping: The system must incrementally construct and maintain semantic maps in real-time, allowing it to adapt to ongoing changes in the environment; 3) Navigation with dynamic changes: In human-centered environments, objects frequently change their location or state—for instance, a laptop might be moved or unfolded. The system should implement a navigation strategy that guides the agent to the relocated object and updates the map for lifelong navigation.
@@ II-A Open-Vocabulary Online Semantic Mapping
Fast closed-set detectors like YOLO [14] have been widely used in semantic robotic mapping [15, 16, 17, 1], but their fixed category sets limit generalization to open-world settings. To enable open-vocabulary perception, two main approaches have emerged. One uses image tagging models [18, 19] to generate class labels from incoming frames, which then guide closed-set detectors [20] for object detection. ConceptGraphs [4] follows this pipeline but remains offline due to the high cost of running GroundedSAM [21] per frame. A more common alternative applies class-agnostic segmentation [7, 8] directly, followed by feature embedding using vision foundation models (VFMs) [22, 23, 24]. HOV-SG [5] uses this method and merges over-segmented objects at the 3D level, improving semantic accuracy but remaining expensive in computation. CLIO [25] addresses efficiency by merging only task-relevant objects. However, since the map is task-specific, querying non-task objects requires extra map reconstruction, limiting map reusability. Different from prior approaches, DualMap adopts a hybrid segmentation strategy for efficient, holistic open-vocabulary mapping and replaces costly 3D inter-object merging with lightweight intra-object checks, enabling online operation.
TABLE I: Comparison of recent semantic mapping systems.
System Open-Vocab Online Mapping Dynamic Handling
Hydra [1]
Khronos [2]
ConceptGraphs [4]
HOV-SG [5]
CLIO [25]
OpenIN [9]
DualMap (Ours)
@@ III-A4 Scene Pointcloud Generation
To support layout estimation for downstream navigation, we maintain a scene point cloud $\mathbf{P}_{\text{s}}$ that captures the overall geometry information of the environment during operation (Fig. 3-c). It is built by selecting RGBD frames with significant pose changes (translation >> 1.0m or rotation >> 20°), projecting their point clouds, and applying downsampling. This process (blue branch in Fig. 3) runs in a parallel low-frequency thread, providing a compact structural representation of the scene.
Figure 3: Observation generation pipeline. From each posed RGBD frame, the system runs three parallel threads: a) Closed-set detection (e.g. YOLO) identifies in-list objects in object-level but misses unseen ones (e.g., stools); b) Open-set segmentation (e.g. FastSAM) captures a broader object range but may over-segment (e.g., window). After refinement and merging, the results are embedded into the final observation set. c) Scene cloud generation runs at low frequency and generates 3D point clouds for layout calculation.
@@ IV-A Map Abstraction
The abstraction of the concrete map involves two steps: first, identifying which objects are selected as anchor objects; and second, associating the remaining volatile objects with anchor objects based on spatial relations. Anchor objects refer to static items that rarely move, such as beds or desks, while volatile objects (e.g., books or cups) frequently change positions. Only anchor objects are retained in the abstract map. We first classify objects using two lists of representative anchor and volatile categories. For each object $o\in\mathcal{M}_{c}$, we compute its similarity to both lists using CLIP features. If the maximum similarity to one list exceeds the other by a margin of $\Delta\tau=0.05$, the object is assigned to the corresponding category. For cases without a clear margin, we further compute the similarity between the object’s feature $\mathbf{f}_{o}$ and the anchor template feature $\mathbf{f}_{\text{t}}$, encoded from descriptive phrases (e.g., “furniture that is not often moved”). If the similarity exceeds a threshold $\tau_{a}$, the object is classified as an anchor object $a$; otherwise, it is a volatile $v$. The threshold $\tau_{a}$ is set to 0.5 based on classification accuracy experiment.
@@
Scene layout, important for navigation, is also extracted in map abstraction. We project the scene point cloud $\mathbf{P}_{\text{s}}$ onto a bird’s-eye-view plane and discretize it into spatial bins. A binary occupancy map is then generated by computing a histogram of point counts across bins and using the 90th percentile as a threshold, highlighting scene structures like walls. The layout is visualized as gray point clouds in Fig. 1, and is updated only when a new navigation query is issued.
@@ IV-C Navigation Strategy
A global path toward $a^{*}$} is planned using a Voronoi-based planner over the abstract map [31]. As the agent moves along this path, it incrementally builds a local concrete map $\mathcal{M}_{c}^{\text{local}}$ and evaluates objects in $\mathcal{M}_{c}^{\text{local}}$ for potential matches. For each object $o\in\mathcal{M}_{c}^{\text{local}}$, we compute the cosine similarity $s(o)$ between its feature $\mathbf{f}_{o}$ and the query feature $\mathbf{f}_{q}$. If $s(o)$ is within a margin $\epsilon$ of $s(a^{*})$ and $o$ lies within the projected footprint of $a^{*}$, it is considered a confident match. A local path is then planned via RRT* [32] to reach the target.
@@ V Experiments
This section first describes the experimental setup and further evaluates DualMap across four core aspects: 1) the effectiveness and efficiency of online concrete map construction; 2) the navigation performance in both static and dynamic scenes; 3) the contribution of key system components (via ablations); and 4) real-world applicability.
@@ V-A1 Datasets
We evaluate DualMap on semantic segmentation, efficiency, and object navigation. For semantic segmentation and efficiency, we use the ScanNet [12] and Replica [11] datasets, following scene selections from prior work [4, 5], with all experiments conducted at the original image resolutions. For static object navigation, we evaluate on the HM3D dataset [13], selecting three representative scenes (00829, 00848, 00880) and augmenting object diversity by integrating YCB objects [33]. For dynamic navigation, as current benchmarks lack object-level dynamics, we construct new environments for dynamic object navigation using a custom tool built on Habitat Simulator [34], which enables dynamic placement of YCB objects. We define two types of object changes: In-anchor relocation, where an object moves within the same anchor (e.g., moving a cup on a table); Cross-anchor relocation, where an object is moved to a different anchor (e.g., from a table to a shelf). For each static HM3D scene, we create dynamic variants by manually repositioning added objects to reflect both types of changes. Agents are deployed in the Habitat Simulator for both static and dynamic navigation experiments. For real-world experiments, we design a rigid perception module with a LiDAR and an RGBD camera, where LiDAR provides pose estimation via FastLIO2 [35]. This module is deployed on both wheeled and quadruped platforms, as shown in Fig. 6-a.
@@
TABLE II: Open-vocabulary 3D Semantic Segmentation and Efficiency
Dataset Method mIoU $\uparrow$ FmIoU $\uparrow$ mAcc $\uparrow$ ODR $\approx 1$ Avg. Mem (MB) $\downarrow$ Peak Mem (MB) $\downarrow$ TPF (s) $\downarrow$
Replica ConceptGraphs 0.1501 0.3858 0.3559 2.02 7148.9 23551.9 4.188
@@
ScanNet ConceptGraphs 0.0882 0.3077 0.3538 6.97 9780.3 26155.2 6.301
@@
V-A2 Evaluation Metrics
For semantic segmentation, we adopt standard metrics from [5], including mean IoU (mIoU), frequency-weighted IoU (F-mIoU), and mean accuracy (mAcc). In both qualitative and quantitative evaluations, we exclude background classes like ["wall", "floor", "ceiling"], which occupy large portions of the scene but provide little semantic information. We also report the Object Density Ratio (ODR), a novel metric defined as the ratio of predicted object counts to ground-truth object counts. This metric captures how well the system reproduces realistic object quantities; values closer to 1 indicate better alignment with real-world object density. ODR complements existing segmentation metrics by revealing over- or under-segmentation issues that may be hidden under high pixel-wise accuracy. For efficiency, we measure average memory usage, peak memory consumption, and time per frame (TPF) during operation. For navigation, we report Success Rate (SR), defined as the percentage of queries in which the agent stops within 1 meter of the queried object. In dynamic scenes, success further requires finding the target within three attempts. Unless otherwise specified, all reported metric values in tables are averaged across scenes for each dataset.
@@ V-A4 Implementation Details
All experiments are conducted on a machine with an NVIDIA RTX 4090 GPU and an Intel i7-12700KF CPU. We adopt the YOLOv8l-world model [20] for closed-set detection, using a class list generated by GPT-4 [30], and supplement it with FastSAM-s [8] to enable open-vocabulary segmentation. For feature embedding, we use MobileCLIP-S2 [24], shared across all methods for fair comparison.
@@ V-B Concrete Mapping Evaluation
We present quantitative and qualitative results in Table II and Fig. 5. DualMap achieves the best semantic segmentation results across datasets, improving mIoU on Replica by 2.8% over HOV-SG and 10.3% over ConceptGraphs. This gain stems from object-level splitting that preserves the diversity of objects and text-enhanced feature embedding. Fig. 5 further illustrates these improvements, with red arrows showing misclassified objects that are better handled by DualMap. For ODR, DualMap achieves the closest alignment with ground truth (0.97 on Replica), thanks to holistic segmentation that reduces over-segmentation and our stability check that removes noisy objects. While HOV-SG and ConceptGraphs apply inter-object merging to control segment counts, they still suffer from redundant fragments. This difference is clearly illustrated in Fig. 5, where noisy RGBD inputs from real sensors result in meaningless predictions (blue arrows) in the baselines. DualMap filters out such errors via stability check, demonstrating robustness in real-world scenarios.
@@
Figure 6: Real-world navigation in dynamic environments. a) Robotic platforms equipped with a perception module that integrates a LiDAR and an RGB-D camera, both mounted on a rigid 3D-printed mount. b–c) Two examples of language-guided navigation in dynamic real-world scenes, where the agent tries to locate relocated objects across multiple attempts.
@@ V-C Navigation in Simulated Environments
TABLE III: Object Navigation Success Rate (SR) on HM3D.
@@
For static navigation, we evaluate all methods across three HM3D scenes. As shown in Table III-Static, DualMap achieves the highest success rate. ConceptGraphs performs competitively due to its object-centric mapping. In contrast, HOV-SG’s segment-merging strategy still results in fragmented maps, leading to more query and navigation failures.
For dynamic scenes (Table III-Dynamic), DualMap naturally handles in-anchor changes, as the queried object remains within its original anchor and can be accurately localized through the local concrete map. For more challenging cross-anchor cases, it updates the abstract map online and reselects reasonable candidates accordingly. This strategy maintains high success rates even under large object movements. Among failure cases in cross-anchor setting, 28.3% are due to false matches, 9.4% to exceeding navigation attempt limits, and 1.9% to planning errors.
@@ V-D1 Mapping Performance
TABLE IV: Semantic Mapping Ablation on Replica
@@
Full System 0.5207 0.4024 0.2538
In Table IV and Fig. 5, we ablate key components of the concrete mapping pipeline on Replica and ScanNet. We first remove FastSAM and rely only on YOLO for detection. Since YOLO detects only predefined categories, many unseen objects are missed, leading to a drop in performance. Next, we disable the refinement step that merges overlapping YOLO detections based on color histogram similarity. Without merging overlaps, objects are over-segmented. The performance drops slightly, as YOLO produces relatively holistic detections. Then, we remove the weighted feature merging and use only image features. This causes a notable drop, indicating the importance of incorporating global semantic cues from text. Finally, we disable the object-level splitting mechanism. This causes under-segmentation, where nearby objects are grouped into one, leading to inaccurate matching and reduced mapping precision. Qualitative results in Fig. 5 show that without object stability check, sensor noise introduces noisy segments, degrading overall semantic quality.
@@ V-D2 Efficiency
TABLE V: Mapping Efficiency Ablation on ScanNet. Gray row indicates configuration used in our proposed system.
@@
In Table V, we analyze three key modules that are related to our mapping pipeline to evaluate their impact on efficiency: Inter-object merging in 3D, intra-object level stability check, and closed-set detection assistance in open-vocabulary segmentation. HOV-SG follows the configuration in Row 1, resulting in the highest time cost and ODR. This is because SAM-style segmentation produces a large number of segments, making 3D-level merging computationally expensive. Configurations in Row 2 improve speed by using YOLO to assist segmentation, but still suffer from noisy objects caused by sensor noise. In Row 3, enabling our proposed stability check significantly improves ODR by removing noisy objects, which also reduces the object merging time cost. Our final configuration (Row 4) disables inter-object merging, slightly sacrificing map accuracy but achieving notable efficiency gains—an important trade-off for real-world robotic applications.
@@ V-D3 Dynamic Scene Navigation
TABLE VI: Success Rates under Different Candidate Selection Strategies for Relocated Objects on HM3D
@@
To evaluate navigation under dynamic changes, we focus on the more challenging cross-anchor setting and test different candidate selection strategies when the queried object has changed position. As shown in Table VI, random selection often fails due to excessive retries. Using the original abstract map $\mathcal{M}_{a}$ provides anchor-based guidance and improves success. Our final strategy updates $\mathcal{M}_{a}$ during navigation to form $\mathcal{M}_{a}^{\prime}$, which makes full use of past navigation information. This helps the agent find the relocated object more effectively.